# FLARE: Faithful Logic-Aided Reasoning and Exploration

### v2

@@ Abstract
Modern Question Answering (QA) and Reasoning approaches based on Large Language Models (LLMs) commonly use prompting techniques, such as Chain-of-Thought (CoT), assuming the resulting generation will have a more granular exploration and reasoning over the question space and scope. However, such methods struggle with generating outputs that are faithful to the intermediate chain of reasoning produced by the model. On the other end of the spectrum, neuro-symbolic methods such as Faithful CoT (F-CoT) propose to combine LLMs with external symbolic solvers. While such approaches boast a high degree of faithfulness, they usually require a model trained for code generation and struggle with tasks that are ambiguous or hard to formalise strictly. We introduce Faithful Logic-Aided Reasoning and Exploration (FLARE), a novel interpretable approach for traversing the problem space using task decompositions. We use the LLM to plan a solution, soft-formalise the query into facts and predicates using a logic programming code and simulate that code execution using an exhaustive multi-hop search over the defined space. Our method allows us to compute the faithfulness of the reasoning process w.r.t. the generated code and analyse the steps of the multi-hop search without relying on external solvers. Our methods achieve SOTA results on $\mathbf{7}$ out of $\mathbf{9}$ diverse reasoning benchmarks. We also show that model faithfulness positively correlates with overall performance and further demonstrate that FLARE allows pinpointing the decisive factors sufficient for and leading to the correct answer with optimal reasoning during the multi-hop search.
@@ 1 Introduction
Complex Reasoning in natural Question Answering (QA) tasks assumes the capability to explore the problem space of the designated query with a formalised set of facts, relations, commonsense knowledge and logical implications. In line with this, LLMs have been enhanced with CoT (Wei et al., 2022) prompting, which supplements the QA process by generating intermediate reasoning chains given a set of in-context examples (Brown et al., 2020a), as shown in Figure 1. This allowed for advancement in commonsense (Madaan et al., 2022), symbolic (Wang et al., 2022; Sprague et al., 2024) and mathematical (Jie et al., 2023) reasoning. Although CoT allows for a problem exploration in natural language steps, such an approach has been shown to cause performance degradation for nuanced reasoning tasks involving multi-hop planning (Valmeekam et al., 2022; Suzgun et al., 2023), problem exploration (Yao et al., 2022) and arithmetic (Hendrycks et al., 2021b; Madaan & Yazdanbakhsh, 2022b). These discrepancies arise as CoT suffers from a limited ability to decompose, search, verify and backtrack using intermediate rationale chains (Yao et al., 2022), cascading hallucinations and errors (Ling et al., 2023) and that natural language might not be an optimal representation for describing the reasoning process (Li et al., 2024). Simultaneously, LLM output has been shown to be unfaithful and inconsistent w.r.t. the intermediate CoT rationale (Jacovi et al., 2024; Lanham et al., 2023b; Turpin et al., 2023).
Figure 1: A depiction of the plan, code and simulated search in FLARE. Each module has a breakdown of the relevant components composed by the LLM explained in Section 2.
@@
To circumvent the problem of CoT faithfulness and allow for more robust reasoning during QA, Xu et al. (2024, Faithful CoT) suggested generating code which is further executed using an external symbolic solver. Producing and executing code allows one to both create an output that can be conditioned on heuristic search in the problem space and perform backtracking. However, strict translations of natural language queries into code, such as autoformalisation (Szegedy, 2020; Wang et al., 2018), is a non-trivial task involving direct inference of implicit commonsense and domain-specific knowledge and the ability to align abstract and informal concepts directly to constrained formal definitions for further execution (Wu et al., 2022). An example query “Do all parts of the aloe vera plant taste good?” seen in Figure 1 is not trivial to formalise, nor to provide a strict algorithmic solution for it, making it much more suitable for soft reasoning. Using external solvers makes such fuzzy reasoning impossible and requires consistently generating syntactically correct executable code. While some LLMs have coding capabilities stemming from their pretraining (Jiang et al., 2024; Aryabumi et al., 2024), relative code consistency is more probable with models explicitly trained for coding (Chen et al., 2021).
@@
To overcome these problems with CoT and F-CoT, we propose Faithful Logic-Aided Reasoning and Exploration (FLARE), an interpretable method that allows for planning, fuzzy reasoning, and traversing the problem space with backtracking, exact task decomposition, and measuring faithfulness. In FLARE, given a natural language query, we prompt an LLM to sequentially generate a plan that includes an analysis and the logical steps necessary for formalising and answering the question, a logical programming Prolog (Wielemaker et al., 2012) code that allows formalising the query into a set of facts, relations and their composition forming the problem space and the search which is an LLM generated simulation of exhaustive multi-hop traversal of that space from the code. An illustration of FLARE can be seen in Figure 1. In our framework, the generated code must not be consistently executable by an external solver, allowing for the soft-formalisation of natural language. Although we see that even generalist LLMs are able to produce executable code in $\geq 50\%$ of cases. FLARE allows us to measure the faithfulness of the outcome w.r.t. the simulated multi-hop logical traversal by directly comparing the search paths generated from executable Prolog code to that LLM generation. This comparison also allows for pinpointing model hallucinations and inconsistencies. We systematically study the effectiveness of our method using $4$ general-purpose LLMs of varying scales across $9$ diverse benchmarks, covering Math World Problems, Multi-hop QA and Relation inference and show that our method achieves state-of-the-art results in $7$ out of $9$ in comparison to CoT and F-CoT. We also show that the method is rather competitive for models tuned for coding, with an average overall increase of $16\%$ over F-Cot and $9\%$ over CoT. Our findings show that model accuracy strongly correlates with the faithfulness of the reasoning process towards the problem space. We also provide ablations showing that the model can interpretably pinpoint hallucinations, search degeneracies, and the limitations of the search over the problem space. Our key contributions are the following:
@@
  • We introduce FLARE a novel paradigm for logic-aided and interpretable formalisation and search over the problem space in QA and reasoning tasks.
@@
  • We perform a systematic evaluation across $9$ benchmarks and $4$ models of varying scales, showing the advantages of using FLARE for QA in a few-shot setup over prior approaches.
@@
  • The modularity of FLARE allows defining a simple ingrained method for measuring model faithfulness and shows that it is strongly correlated with performance.
@@
2 Faithful Logic-Aided Reasoning and Exploration
@@ 2.1 LLM Simulated Search
FLARE comprises three modules for generating a plan, code and simulated search for answering a natural language query $\mathcal{Q}=\{T_{1}\dots T_{|\mathcal{Q}|}\}$, where each $T_{i}$ is a token.
@@
For each query $\mathcal{Q}$, given an LLM $\mathcal{M}$, we initially prompt it to generate a plan, $\mathcal{P}$, which should be comprised of task explanation, analysis and a plan for further formalising the query. An example of this can be seen in the plan section in Figure 1. We use in-context few shot examples $\mathcal{E}_{\mathcal{P}}$ of such plan generations along with greedy decoding for obtaining the final plan.
$\mathcal{P}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T_{i}\mid\mathcal{E}_{\mathcal{P}},\mathcal{Q})$ (1)
@@
After generating the plan, we prompt the LLM $\mathcal{M}$ to generate a Prolog code $\mathcal{C}$, an example of which can be seen in Figure 1. We append executable code generation samples $\mathcal{C}_{\text{sample}}$ to the previous in-context examples $\mathcal{E}_{\mathcal{P}}$ and obtain few-shot code generation demonstrations $\mathcal{E}_{\mathcal{C}}=[\mathcal{E}_{\mathcal{P}};\mathcal{C}_{\text{sample}}]$
$\mathcal{C}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T_{i}\mid\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{P})$ (2)
@@
Prolog is a symbolic logic-programming engine (Bowen, 1979) used for heuristic search over Horn Clauses (Chandra & Harel, 1985). It is a declarative programming paradigm (Lloyd, 1994), meaning that the code is expressed as the logic of computation. In particular, this logic is formalised as a set of facts $\mathcal{F}$ and relations $\mathcal{R}$ forming our problem space, while the final goal $\mathcal{G}$ is a first-order logic combination of them. As a default, Prolog uses a depth-first search (DFS) strategy (Bowen, 1979) for sub-goal decomposition and feasible traversal of the problem space that satisfies the goal $\mathcal{G}$. Such a traversal is referred to as the trace. At each trace step, the program can either confirm or invalidate the sub-goal using the feasibility of fact and relation combinations, expand the search tree or retry satisfying a failed sub-goal with new combinations. An example of such a search can be observed in Figure 1. It is possible to complete an exhaustive search, exploring all possible paths that do or do not satisfy the goal. These characteristics are beneficial as we can explicitly access and segment the facts and relations that form the problem space and the search strategy used for query formalisation. As Prolog is declarative, it is sufficient to use a regexp heuristic for the segmentation, which is referred to as EXTRACT in Equation 2 and Equation 3. Furthermore, including exhaustive traversal traces in-context allows the LLM to simulate sub-goal decomposition, backtracking, intermediate goal invalidation, etc. We discuss this in more depth in the next paragraph.
@@
After generating the logic-programming code, we want to simulate program execution by generating a problem space traversal trace with our LLM $\mathcal{M}$. We update our in-context samples by appending search traces $\mathcal{S}_{\text{sample}}$ constructed from Prolog execution of sample codes $\mathcal{C}_{\text{sample}}$, i.e. $\mathcal{E}_{\mathcal{S}}=[\mathcal{E}_{\mathcal{C}};\mathcal{S}_{\text{sample}}]$.
$\mathcal{S}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T_{i}\mid\mathcal{E}_{\mathcal{S}},\mathcal{Q},\mathcal{P},\mathcal{C})$ (3)
@@
During iterative problem space traversal, we can segment the facts $\mathcal{F}_{\text{\emph{search}}}$, relations $\mathcal{R}_{\text{\emph{search}}}$, completed and backtracked paths with their answers $\mathcal{A}_{\text{\emph{search}}}$ used during the search simulation. To get the final answer we update in-context samples with the correct final answers $\mathcal{A}_{\text{sample}}$ from the executed search $\mathcal{S}_{\text{sample}}$, $\mathcal{E}_{\mathcal{A}}=[\mathcal{E}_{\mathcal{S}};\mathcal{A}_{\text{sample}}]$.
$\mathcal{A}_{\text{\emph{Final}}}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T_{i}\mid\mathcal{E}_{\mathcal{A}},\mathcal{Q},\mathcal{P},\mathcal{C},\mathcal{})$ (4)
The prompts used for generating each part in FLARE can be seen in Appendix A.

### v3

@@ Abstract
Modern Question Answering (QA) and Reasoning approaches based on Large Language Models (LLMs) commonly use prompting techniques, such as Chain-of-Thought (CoT), assuming the resulting generation will have a more granular exploration and reasoning over the question space and scope. However, such methods struggle with generating outputs that are faithful to the intermediate chain of reasoning produced by the model. On the other end of the spectrum, neuro-symbolic methods such as Faithful CoT (F-CoT) and Logic-LM propose to combine LLMs with external symbolic solvers. While such approaches boast a high degree of faithfulness, they usually require a model trained for code generation and struggle with tasks that are ambiguous or hard to formalise strictly. We introduce Faithful Logic-Aided Reasoning and Exploration (FLARE), a novel interpretable approach for traversing the problem space using task decompositions. We use the LLM to plan a solution, formalise the query into facts and predicates, which form the problem space, using a logic programming code and simulate that code execution using an exhaustive multi-hop search over the defined space. Our method allows us to compute the faithfulness of the reasoning process w.r.t. the generated code and explicitly trace the steps of the multi-hop search without relying on external solvers. Our methods achieve SOTA results on $\mathbf{7}$ out of $\mathbf{9}$ diverse reasoning benchmarks. We also show that model faithfulness positively correlates with overall performance and further demonstrate that FLARE allows pinpointing the decisive factors sufficient for and leading to the correct answer with optimal reasoning during the multi-hop search. Our findings reveal that successful traces exhibit, on average, a $18.1\%$ increase in unique emergent facts, a $8.6\%$ higher overlap between code-defined and execution-trace relations, and a $3.6\%$ reduction in unused code relations.
@@ 1 Introduction
Complex Reasoning in natural Question Answering (QA) tasks assumes the capability to explore the problem space of the designated query with a formalised set of facts, relations, commonsense knowledge and logical implications. In line with this, LLMs have been enhanced with CoT (Wei et al., 2022) prompting, which supplements the QA process by generating intermediate reasoning chains given a set of in-context examples (Brown et al., 2020a), as shown in Figure 1. This allowed for advancement in commonsense (Madaan et al., 2022), symbolic (Wang et al., 2022; Sprague et al., 2024) and mathematical (Jie et al., 2023) reasoning. Although CoT allows for a problem exploration in natural language steps, such an approach has been shown to cause performance degradation for reasoning tasks involving multi-step planning (Valmeekam et al., 2022; Suzgun et al., 2023), problem exploration (Yao et al., 2022), and arithmetic tasks (Hendrycks et al., 2021b; Madaan & Yazdanbakhsh, 2022b). These discrepancies arise as CoT suffers from a limited ability to decompose, search, verify and backtrack using intermediate rationale chains (Yao et al., 2022), cascading hallucinations and errors (Ling et al., 2023) and that natural language might not be an optimal representation for describing the reasoning process (Li et al., 2024). Simultaneously, LLM output has been shown to be unfaithful and inconsistent w.r.t. the intermediate CoT rationale (Jacovi et al., 2024; Lanham et al., 2023b; Turpin et al., 2023).
Figure 1: A depiction of the plan, code and simulated search in FLARE. Each module is generated separately and iteratively, allowing us to obtain the final answer. The green and yellow highlighted text shows the overlap between the facts and the relations between the code and the simulated search.
@@
To mitigate the problem of CoT faithfulness and allow for more robust reasoning during QA, Lyu et al. (2023, Faithful CoT) and Logic-LM (Pan et al., 2023) suggested generating code which is further executed using an external symbolic solver. Producing and executing code enables the generation of outputs guided by external solvers, leveraging search with backtracking to explore the problem space effectively. However, strict translations of natural language queries into code, such as autoformalisation (Szegedy, 2020; Wang et al., 2018), is a non-trivial task involving direct inference of implicit commonsense and domain-specific knowledge and the ability to align abstract and informal concepts directly to constrained formal definitions for further execution (Wu et al., 2022). An example query, “Do all parts of the aloe vera plant taste good?”, is challenging to formalize or address with a strict algorithmic solution, as it requires interpretative, deductive and context-dependent reasoning, referred to as soft or fuzzy reasoning. Using external solvers makes such fuzzy reasoning impossible and requires consistently generating syntactically correct executable code. While some LLMs have coding capabilities stemming from their pretraining (Jiang et al., 2024; Aryabumi et al., 2024), relative code consistency is more probable with models explicitly trained for coding (Chen et al., 2021).
@@
To overcome these problems, we propose Faithful Logic-Aided Reasoning and Exploration (FLARE), an interpretable method that allows for planning, fuzzy reasoning, and traversing the problem space with backtracking, exact task decomposition, and measuring faithfulness. In FLARE, given a natural language query, we prompt an LLM to sequentially generate a plan that includes an analysis and the logical steps necessary for formalising and answering the question, a logic programming (Wielemaker et al., 2012) code that allows formalising the query into a set of facts, relations and their composition forming the space for exploring that query and the search, which is an LLM-generated code execution simulation. An illustration of FLARE can be seen in Figure 1. In our framework, the generated code must not be consistently executable by an external solver, allowing for the soft-formalisation of natural language. Although we see that even generalist LLMs are able to produce executable code in $\geq 50\%$ of cases. FLARE allows us to measure the faithfulness of the outcome w.r.t. the simulated code execution by directly comparing the search paths produced by the external solver to that LLM generation. This comparison also allows for pinpointing model hallucinations and inconsistencies. We systematically study the effectiveness of our method using $4$ general-purpose LLMs of varying scales across $9$ diverse QA and $3$ logical inference benchmarks, covering Math World Problems, Multi-hop QA, Relation inference, deductive and analytical reasoning and show that our method achieves state-of-the-art results in $7$ out of $9$ QA datasets and $2$ out of $3$ logic datasets in comparison to CoT, F-CoT and Logic-LM. We also show that the method is competitive for models tuned for coding, with an average overall increase of $16\%$ over F-Cot and $9\%$ over CoT. Our findings show that model accuracy strongly correlates with the faithfulness of the reasoning process towards search traces from the simulated code execution. We also provide ablations showing that the model can interpretably pinpoint hallucinations, underutilized knowledge, and the limitations of the search over the problem space. Our key contributions are the following:
@@
  • We introduce FLARE a novel paradigm for logic-aided and interpretable formalisation and search over the problem space in QA and logic reasoning tasks.
@@
  • We perform a systematic evaluation across $9$ QA and $3$ logical inference benchmarks and $4$ models of varying scales, showing the advantages of using FLARE for QA in a few-shot setup over prior approaches.
@@
  • The modularity of FLARE allows defining a simple ingrained method for measuring model faithfulness, which is further shown to be strongly correlated with performance.
@@
2 Related Work
Reasoning in Natural Language
Reasoning with Search
Several lines of work propose using techniques to expand the reasoning paths over the problem space. Self-consistency decoding (Wang et al., 2023b) is an approach used to sample many natural language reasoning paths and take a majority vote for an answer. Another popular approach is Tree-of-Thoughts (ToT; Yao et al., 2023a), which proposes to explore the query with reasoning similar to a tree traversal, where each state is created and evaluated using an LLM. Similar techniques try to adapt symbolic search approaches akin to DFS, BFS (Besta et al., 2024), A (Lehnert et al., 2024) or other combinations (Gandhi et al., 2024) with direct tuning (Lehnert et al., 2024), imitation training (Yang et al., 2022) or few-shot prompting (Zhang et al., 2024). It must be noted that all of these techniques have only been tested in constrained mathematical puzzle-solving and algorithmic domains like the 24 Game (Yang et al., 2022), Countdown (Wikipedia, 2024), Sorting (Besta et al., 2024), maze solving (Yang et al., 2022), Sokoban (Lehnert et al., 2024), and others. Although the search component of FLARE has some similarities to these techniques, we argue that our method allows for generalistic reasoning with interpretable multi-hop search through simulated code execution.
Reasoning with Formalisation
Another line of research has tried formalising natural language queries into code (Gao et al., 2023; Li et al., 2024) or pseudo-code (Chae et al., 2024; Gandhi et al., 2024). This allows the translation of the query into a strict structure and offloads the reasoning and search components to deterministic solvers like Python Chen et al. (2023), PDDL Lyu et al. (2023); Liu et al. (2023), DataLog Lyu et al. (2023) and others. While models are capable of synthesising programs (Austin et al., 2021; Nijkamp et al., 2023) and benefit from the use of code in numerical and algorithmic reasoning settings (Chen et al., 2023; Gao et al., 2023), the usage of code for general QA has not been rigorously explored. The reasons are that formalisation from natural language into a strict and executable code is challenging (Wu et al., 2022), following the exact syntactic constraints of the programming language not abundantly used during pre-training is onerous (Liu et al., 2024) and can require models explicitly tuned for coding (Chen et al., 2021). Using an external solver for reasoning also limits the capability for soft reasoning in commonsense knowledge and implications. Although we formalise the natural language query into a logic programming Prolog program during the code generation part of FLARE, we do not explicitly require the code to be executable and do not use external solvers during inference. This allows for the further use of the LLM for soft-reasoning to simulate code execution in a logic-based problem space traversal similar to Prolog while circumventing the need for code tuning a generalist model.
[…]
Reasoning Faithfulness
3 Methodology
@@ 3.1 LLM Simulated Search
FLARE comprises three modules for generating a plan, code and simulated search for answering a natural language query $\mathcal{Q}=\{T^{\mathcal{Q}}_{1}\dots T^{\mathcal{Q}}_{|\mathcal{Q}|}\}$, where each $T^{\mathcal{Q}}_{i}$ is a token in the query $\mathcal{Q}$.
@@
For each query $\mathcal{Q}$, given an LLM $\mathcal{M}$, we initially use instructions $\mathcal{I}^{\mathcal{P}}$ to prompt it to generate a plan $\mathcal{P}$, which should be comprised of task explanation, analysis and a plan for further formalising the query. An example of this can be seen in the plan section in Figure 1. We use in-context few shot examples $\mathcal{E}_{\mathcal{P}}$ of such plan generations along with greedy decoding for obtaining the final plan.
$\mathcal{P}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T^{\mathcal{P}}_{i}\mid T^{\mathcal{P}}_{:i-1},\mathcal{E}_{\mathcal{P}},\mathcal{Q},\mathcal{I}^{\mathcal{P}})$ (1)
Where $\mathcal{P}_{i}\text{ and }T^{\mathcal{P}}_{i}$ is the $i$-th token in the generated plan $\mathcal{P}$ and $p_{\mathcal{M}}$ is the probability of the next token over the vocabulary obtained from model $\mathcal{M}$.
@@
After generating the plan, we use instructions $\mathcal{I}^{\mathcal{C}}$ to prompt the LLM $\mathcal{M}$ to generate a Prolog code $\mathcal{C}$, an example of which can be seen in Figure 1. We append executable code generation samples $\mathcal{C}_{\text{sample}}$ to the previous in-context examples $\mathcal{E}_{\mathcal{P}}$ and obtain few-shot code generation demonstrations $\mathcal{E}_{\mathcal{C}}=[\mathcal{E}_{\mathcal{P}};\mathcal{C}_{\text{sample}}]$
$\mathcal{C}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T^{\mathcal{C}}_{i}\mid T^{\mathcal{C}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}})$ (2)
@@
Where $\mathcal{C}_{i}\text{ and }T^{\mathcal{C}}_{i}$ is the $i$-th token in the generated code $\mathcal{C}$.
Figure 2: The trend of mean model accuracy w.r.t mean faithfulness for all the models.
@@
Prolog is a symbolic logic-programming engine (Bowen, 1979) designed for heuristic search over Horn Clauses (Chandra & Harel, 1985). As a declarative programming paradigm (Lloyd, 1994), the code is expressed as the logic of computation, expressed as a set of facts $\mathcal{F}$ and relations $\mathcal{R}$ defining the problem space, with the goal $\mathcal{G}$ being a first-order logic combination of them. Prolog employs depth-first search (DFS) (Bowen, 1979) for sub-goal decomposition and traversal of the problem space, satisfying $\mathcal{G}$ through a sequence of steps known as the trace. Each step either confirms/invalidates a sub-goal, expands the search tree, or retries failed sub-goals with new combinations. An example of such a search is shown in Figure 1. Prolog supports exhaustive search by exploring all paths that satisfy or fail the goal. This explicit segmentation of facts, relations, and search strategies simplifies query formalization. As a declarative language, Prolog enables segmentation using a simple regexp heuristic, referred to as EXTRACT in Equation 2 and Equation 3. Including exhaustive traces in-context allows an LLM to simulate sub-goal decomposition, backtracking, and intermediate goal invalidation, discussed further in the next paragraph.
@@
After generating the logic-programming code, we want to simulate program execution by generating a problem space traversal trace with our LLM $\mathcal{M}$. We use instructions $\mathcal{I}^{\mathcal{S}}$ and update our in-context samples by appending search traces $\mathcal{S}_{\text{sample}}$ constructed from Prolog execution of sample codes $\mathcal{C}_{\text{sample}}$, i.e. $\mathcal{E}_{\mathcal{S}}=[\mathcal{E}_{\mathcal{C}};\mathcal{S}_{\text{sample}}]$.
$\mathcal{S}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T^{\mathcal{S}}_{i}\mid T^{\mathcal{S}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}},\mathcal{C},\mathcal{I}^{\mathcal{S}})$ (3)
@@
Where $T^{\mathcal{S}}_{i}$ is the $i$-th token in the generated search trace $\mathcal{S}$. During iterative problem space traversal, we can segment the facts $\mathcal{F}_{\text{\emph{search}}}$, relations $\mathcal{R}_{\text{\emph{search}}}$, completed and backtracked paths with their answers $\mathcal{A}_{\text{\emph{search}}}$ used during the search simulation. To get the final answer we update in-context samples with their correct final answers $\mathcal{A}_{\text{sample}}$ from the executed search $\mathcal{S}_{\text{sample}}$, $\mathcal{E}_{\mathcal{A}}=[\mathcal{E}_{\mathcal{S}};\mathcal{A}_{\text{sample}}]$ and use instructions $\mathcal{I}^{\mathcal{A}}$ to obtain the final answer from the model.
$\mathcal{A}_{\text{\emph{Final}}}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T^{\mathcal{A}}_{i}\mid T^{\mathcal{A}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}},\mathcal{C},\mathcal{I}^{\mathcal{S}},\mathcal{S},\mathcal{I}^{\mathcal{A}})$ (4)
The prompts used for generating each part in FLARE can be seen in Appendix A along with a complete example in Table 8 and a pseudo-code in Algorithm 1.